OpenStudy (anonymous):

78 4/7 % Find fraction notation and simplify

OpenStudy (vishal_kothari):

5/7....

OpenStudy (vishal_kothari):

78 4/7 = (78*7)/7 + 4/7 = 550/7 Now put that into your fraction: (78 4/7)/100 --> (500/7)/100 Now simplify: 500/7/100 --> (5/7)*(100/100) = 5/7*1 = 5/7..